2A +B 🔛product is an elementary reaction , if pressure is increased three times of the initial pressure, the velocity of forward reaction will be ____ of the previous velocity

Grade:11

2 Answers

Soumodeep
6 years ago
See rate=[P/RT]^2[P/RT]. So for initial rate is [P/RT]^2[P/RT]and new rate is [3P/RT]^2[3P/RT]=27×initial rate so R1/R2=27
Phulkant Kumar
2 years ago
We know rate of reaction is as
R=K(A)x(B)
This law holds for gaseous as well where A and B are pressure of gases and k becomes velocity constant
Now 
we can write this rxn. as  ROR = K(A)2(B)
when pressure is increased three times
 ROR` = (3A)2(3B) 
   SO ROR`= 27(A)2(B)
              ROR` = 27 ROR   ( ROR=A2B)